OpenStudy (anonymous):

I got this question wrong and I do not understand why. lx-3l = 5 x = -8. Since the x is in the absolute value bars, I replaced it with -8 which gave me 8-3 which does equal 5. Where did I go wrong here?

OpenStudy (amistre64):

Is it |x-3| = 5 = -8?

OpenStudy (amistre64):

* 5x of course

OpenStudy (anonymous):

It was a True or False Question. It says "One solution of: lx-3l = 5 is x = -8"

OpenStudy (amistre64):

thats easier to understand now... now show me what you did to get your answer...

OpenStudy (anonymous):

I replaced x with -8 so the problem read l-8-3l = 5. Since the -8-3 was in absolute value bars, I read it as 8-3 which equals 5.

OpenStudy (amistre64):

you did good to begine with; but what the absolute value bars are telling us is the the "equation" inside of them...that value will always become positive. What does -8-3= ? then |?|

OpenStudy (anonymous):

a - minus a + is is actually a negative plus a negative, so it would be -11. Okay, I get that, but what I do not get is I thought absolute value bars mean to ignore any positive or negative signs and read the number for exactly what it's value is UNLESS there is a sign outside of the bars, for example, if it was -l-8-3l.

OpenStudy (amistre64):

your are correct in your understanding of the absolute value bars, but what you are missing is that those bars only apply to the "value" inside of them and not to any of the individual parts that make up that value. since -8-3 = -11 then |-8-3| = |-11| = 11. Does that make sense?

OpenStudy (anonymous):

It does. I did not realize the bars only applied to the value and not to any negative sign. I get it now, thank you very much!

OpenStudy (amistre64):

youre welcome :)